# Jack of All Trades, Master of Some, a Multi-Purpose Transformer Agent

### v1

@@ Abstract
The search for a general model that can operate seamlessly across multiple domains remains a key goal in machine learning research. The prevailing methodology in Reinforcement Learning (RL) typically limits models to a single task within a unimodal framework, a limitation that contrasts with the broader vision of a versatile, multi-domain model. In this paper, we present Jack of All Trades (JAT), a transformer-based model with a unique design optimized for handling sequential decision-making tasks and multimodal data types. The JAT model demonstrates its robust capabilities and versatility by achieving strong performance on very different RL benchmarks, along with promising results on Computer Vision (CV) and Natural Language Processing (NLP) tasks, all using a single set of weights. The JAT model marks a significant step towards more general, cross-domain AI model design, and notably, it is the first model of its kind to be fully open-sourced [^1] https://huggingface.co/jat-project/jat, including a pioneering general-purpose dataset.
@@ 1 Introduction
While combining visual and textual tasks has been well-researched, integrating RL tasks remains relatively unexplored and poses distinct challenges. RL tasks are inherently diverse and heterogeneous, making their combination among themselves and with other domains a highly complex exercise. This integration requires dealing with a landscape of different modalities, task complexities, and data volumes across domains and tasks. New questions that arise include: (1) How to design a model and learning method that effectively handles different modalities and data types (sequential and non-sequential)? (2) How to formulate a learning objective that appropriately balances and harmonizes the different modalities, tasks, and domains without bias toward any particular domain or task? (3) How to design a learning strategy that can accommodate the different levels of complexity inherent in different tasks?
These goals are concurrent and, to our knowledge, have only been addressed together by Reed et al. (2022) with the Gato model. Our contributions are characterized by three major advances: (1) Our model features an innovative structure optimized for sequential tasks. It uniquely assigns each timestep to a corresponding token embedding, resulting in a simpler design. This approach significantly expands the attention window in terms of timesteps compared to Gato (e.g., it is 19 times larger for Atari and more than 25 times larger for Meta-World). (2) In the spirit of open source, we release our code, dataset, and model to the research community. (3) We add observation prediction as an auxiliary task to our model. We demonstrate that this integration significantly contributes to learning a more efficient agent.
@@ 2.1 Transformer for RL
Transformer models (Vaswani et al., 2017) are designed to model sequences and in particular sequences of words in natural language. However, sequence modeling problems span over a much larger set of domains than only NLP. In this paper, we focus on modeling RL trajectories (i.e., sequences of observations, actions and rewards). Modeling such sequences with a Transformer was introduced by Chen et al. (2021) and the Decision Transformer (DT) model. In DT, a Transformer model is trained with offline RL to take sequences of transitions as input and predict the next action. In particular, Chen et al. (2021) proposed to use returns-to-go (i.e. the return from the current state) to condition actions’ generation on both the previous observation and the desired return-to-go. While this has the advantage of explicitly modeling the relations between action selection and return, using the model at inference requires providing at every step a desired return. Liu & Abbeel (2023) proposed to extend this by using hindsight relabelling to better exploit sub-optimal trajectories. Zheng et al. (2022) also extended the DT approach by mixing offine pretraining and online finetuning. Finally, Lee et al. (2022) studied how the DT approach scales to a multi-task RL setup where a single policy is learned for multiple games. Our work lies in this line of work as it also leverages Transformers to model trajectories. However, our approach (1) uses standard Behavior Cloning (BC) instead of conditional BC, relaxing the need to condition the agent by the return-to-go and (2) models a multi-task dataset in which sequences come from very different domains (e.g. control, Atari, visual question answering, see Section 3.2).
@@ 3 Methodology
In this section, we introduce the JAT model, detailing our architectural choices that underpin its effectiveness and highlighting its ability to handle different modalities in both sequential control and text-centric tasks. We present the associated dataset, which is notable for its groundbreaking diversity across domains and modalities. Finally, we discuss in depth the learning strategy used.
@@ 3.1.1 Embedding Mechanism
The model is designed to handle two main categories of data: tasks involving sequential decision-making and non-sequential tasks. In non-sequential tasks, the model currently supports two modalities: text and image. Although the current version of the model supports image generation, we focus only on tasks that involve text generation. To ease reading, we will refer to it as text-centric tasks in the remainder of this paper. Each of these two categories requires a slightly different approach to the embedding process.
@@
For tasks involving sequential decision-making, the data comprises a sequence of observations, actions, and rewards. At the embedding stage, these sequences are processed to produce an interleaved sequence of observation embeddings (augmented with the corresponding reward) and action embeddings, denoted as $[\phi(s_{0},0.0),\phi(a_{0}),\phi(s_{1},r_{1}),\phi(a_{1}),\ldots]$. Unlike DT (Chen et al., 2021) and Gato (Reed et al., 2022), each timestep is consistently associated with two embeddings: one for the observation and the other for the action, regardless of the modality. This enables JAT to better handle high-dimensional observations, and to provide a much wider, constant attention window in terms of timesteps. As an example, this multiplies the size of the attention window in terms of timesteps by more than 25 for Meta-World. The embedding method employed at a specific timestep is modality-dependent (with $H$ the hidden size of the model):
@@ 4 Experiments and Results
In this section we discuss the results of our experiments. First, we provide a brief overview of the model’s performance on text-centric tasks. We then present the results of the sequential decision tasks, showing the different levels of mastery across the different domains within our study. Finally, we provide a comprehensive analysis highlighting the benefits of incorporating the prediction of the next observation as an auxiliary task during the learning process.
@@ 4.2 Sequential Decision-Making Tasks
The final agent achieves a mean expert normalized interquartile mean (IQM) of 63.5%, demonstrating the network’s ability to effectively mimic expert agents across a wide range of tasks. The agent achieves 6.1% of the expert’s score on the Atari 57 benchmark, corresponding to 16.8% of human performance, and exceeding the average human level in 10 games. For the BabyAI benchmark, JAT achieved a normalized score of 98.9%. However, this score falls below 50% for two tasks, namely Boss Level No Unlock and Move Two Across S8N9. For this benchmark, however, there is no guarantee that the expert score can be achieved, since the bot used for dataset collection has access to the full state of the environment, while the interacting agents only have access to a partial observation. Finally, in the MuJoCo and Meta-World, JAT records scores of 81.2% and 68.0%, respectively. Although JAT reaches the expert level for a fair number of Meta-World tasks, we note that a few, such as Basketball, fail completely to be learned. Insofar as the action and observation spaces are identical for all tasks in this benchmark, these failures may be due to task indeterminacy, which we explore in more detail in Appendix D. Future research will have to confirm this hypothesis. We also note that some domains are mastered more quickly than others; in particular, BabyAI achieves a score of 90% after only 30,000 learning steps. We hypothesize that the high semantic similarity of the tasks enables a strong positive transfer, without however providing any proof of this. The Appendix A presents the final results in detail.
@@
Although the results achieved are commendable, for a fair comparison we limit our benchmarking to Gato only, as it is the only truly comparable baseline. Reed et al. (2022) present results only for the 1.18 billion parameter version of Gato, which is 6 times larger than JAT. Its results are normalized to expert performance. Since we don’t have access to the normalization parameters, we estimated scores for the random agents, which may not be exactly the same as those used by Reed et al. (2022), and used our expert scores for normalization, even though they obviously do not match those used by Reed et al. (2022). Therefore, comparisons of these normalized scores should be interpreted with great caution. On the Atari benchmark, JAT achieves an average normalized score of 20.1% approaching Gato, which reports a score of 30.9%. For BabyAI, JAT achieves an average normalized score of 86.0%, also close to the Gato score of 93.2%. Our study, however, is made with 39 tasks versus the 46 used in Gato’s training, with the specific seven additional tasks in their study remaining unidentified. Since our evaluation includes all of the hardest tasks mentioned in their study, the seven missing tasks are likely to be easier, suggesting a harder test scenario in our study. For Meta-World, JAT achieves an average normalized score of 64.5%, which is below the 87.0% reported for Gato. On the MuJoCo benchmark, JAT achieves an average normalized score of 72.5%. While Gato’s training doesn’t use MuJoCo, it’s worth noting that they use the DMC benchmark, which shares some similarities. For reference, on the DMC benchmark (Tassa et al., 2018), Gato achieves an average score of 63.6%.
@@ 5 Conclusion
In this study, we introduce JAT, a novel multimodal framework for general RL agents. JAT features the ability to handle diverse tasks of varying complexity using a single set of parameters. Its innovations include a new transformer-based structure that efficiently addresses sequential control, CV, and NLP tasks. We also show that joint learning of observation prediction significantly improves performance in sequential control tasks. We’ve open-sourced our training dataset, which includes a wide range of sequential control data as well as extensive language and visual data. We believe that JAT represents an important and valuable step towards general-purpose RL models.
@@ Acknowledgements
[…]
@@ Appendix A Full results
This appendix contains a detailed view of the results of the trained JAT agent. The score of the random agent for Atari games is sourced from (Mnih et al., 2015). In other domains, this score is approximated by averaging the returns from 1,000 episodes, where the agent selects actions uniformly across its action space. The expert scores represent the average return in the dataset for the task. Meanwhile, the raw score is the average return achieved by the trained agent, based on 100 evaluation episodes. Both these scores, along with the trained agent, are accessible as open-source [^6] https://huggingface.co/jat-project/jat. The normalized score is derived by comparing the agent’s return to the expert’s, calculated using the formula: $\frac{\text{{score}}-\text{{random\_score}}}{\text{{expert\_score}}-\text{random\_score}}$. It’s important to note that in instances where the expert, inaccurately named, does not fully master the task and thus scores similarly or lower than the random agent, the normalized score must be interpreted cautiously. Specifically, if this score falls below that of the random agent, as in the case of Bowling, normalization is not applied. The results for Atari are presented in Table 1, for BabyAI in Table 2, for Meta-World in Table 3, and for MuJoCo in Table 4.
@@
Alien 227.8 16,912.5 ± 7087.4 1085.9 ± 396.4 0.05 ± 0.02
Amidar 5.8 2164.7 ± 1229.5 41.3 ± 28.6 0.02 ± 0.01
Assault 222.4 15699.1 ± 9572.1 772.9 ± 59.3 0.04 ± 0.00
Asterix 210.0 3699.6 ± 2421.3 778.5 ± 429.0 0.16 ± 0.12
Asteroids 719.0 $1$ 35334.2 1423.6 ± 538.8 0.00 ± 0.00
Atlantis 12850.0 320679.6 ± 418247.4 23541.0 ± 10376.7 0.03 ± 0.03
Bank Heist 14.2 1322.4 ± 60.8 685.5 ± 157.9 0.51 ± 0.12
Battle Zone 236.0 295592.6 ± 161961.0 12950.0 ± 4306.7 0.04 ± 0.01
Beam Rider 363.9 29589.3 ± 16133.0 762.0 ± 243.3 0.01 ± 0.01
Berzerk 123.7 57085.3 ± 13104.5 523.9 ± 161.9 0.01 ± 0.00
Bowling 23.1 20.4 ± 7.3 30.0 ± 11.5 N/A
Boxing 0.1 98.0 ± 3.8 87.0 ± 22.6 0.89 ± 0.23
Breakout 1.7 703.0 ± 203.6 9.2 ± 5.8 0.01 ± 0.01
Centipede 2090.9 11624.3 ± 4918.3 4461.7 ± 2188.8 0.25 ± 0.23
Chopper Command 811.0 90990.6 ± 270876.9 1497.0 ± 723.1 0.01 ± 0.01
Crazy Climber 10780.5 179296.9 ± 39862.1 52850.0 ± 31617.9 0.25 ± 0.19
Defender 2874.5 351958.3 ± 40466.8 10627.5 ± 4473.2 0.02 ± 0.01
Demon Attack 152.1 92195.2 ± 26174.8 315.1 ± 279.0 0.00 ± 0.00
Double Dunk -18.6 20.9 ± 3.6 0.1 ± 11.6 0.47 ± 0.29
Enduro 0.0 $2$ 147.5 111.5 ± 27.4 0.05 ± 0.01
Fishing Derby -91.7 7.2 ± 25.1 -55.2 ± 19.4 0.37 ± 0.20
Freeway 0.0 33.9 ± 0.3 24.1 ± 1.6 0.71 ± 0.05
Frostbite 65.2 13196.1 ± 4341.0 617.3 ± 686.1 0.04 ± 0.05
Gopher 257.6 81676.2 ± 46329.5 2947.2 ± 1448.3 0.03 ± 0.02
Gravitar 173.0 3986.6 ± 1729.0 1030.5 ± 719.2 0.22 ± 0.19
H.E.R.O. 1027.0 44677.4 ± 1754.4 6997.9 ± 2562.5 0.14 ± 0.06
Ice Hockey -11.2 25.2 ± 5.8 -3.8 ± 3.1 0.20 ± 0.09
James Bond 29.0 27786.9 ± 33819.2 187.5 ± 72.2 0.01 ± 0.00
Kangaroo 52.0 574.0 ± 636.9 124.0 ± 156.9 0.14 ± 0.30
Krull 1598.0 11439.8 ± 1218.3 8933.0 ± 1358.6 0.75 ± 0.14
Kung-Fu Master 258.5 32392.8 ± 10006.6 100.0 ± 142.1 -0.00 ± 0.00
@@
Ms. Pacman 307.3 6896.1 ± 2032.0 1516.3 ± 376.7 0.18 ± 0.06
Name This Game 2292.3 22991.2 ± 2473.1 3798.6 ± 1361.6 0.07 ± 0.07
Phoenix 761.5 424583.2 ± 97649.2 1267.5 ± 1013.7 0.00 ± 0.00
PitFall -229.4 -1.4 ± 4.5 -287.4 ± 492.8 -0.25 ± 2.16
Pong -20.7 21.0 ± 0.2 -11.0 ± 11.3 0.23 ± 0.27
Private Eye 24.9 100.0 ± 0.0 96.0 ± 19.6 0.95 ± 0.26
Q*Bert 163.9 42971.4 ± 85070.7 1701.8 ± 1912.6 0.04 ± 0.04
River Raid 1338.5 14800.9 ± 7924.6 2793.1 ± 693.8 0.11 ± 0.05
Road Runner 11.5 77942.8 ± 6088.6 7699.0 ± 3446.6 0.10 ± 0.04
Robotank 2.2 80.5 ± 13.3 16.4 ± 5.2 0.18 ± 0.07
Seaquest 68.4 2597.3 ± 386.1 515.2 ± 141.5 0.18 ± 0.06
Skiing -17098.0 -10738.1 ± $1$29396.1 ± 3289.8 -1.93 ± 0.52
Solaris 1236.3 1353.7 ± 517.0 988.2 ± 487.4 -2.11 ± 4.15
Space Invaders 148.0 29425.3 ± 23623.9 339.5 ± 164.1 0.01 ± 0.01
Star Gunner 664.0 360588.6 ± 49207.7 978.0 ± 638.4 0.00 ± 0.00
Surround -10.0 9.4 ± 0.8 -8.2 ± 1.2 0.09 ± 0.06
Tennis -23.8 11.1 ± 7.6 -22.4 ± 2.2 0.04 ± 0.06
Time Pilot 3568.0 69583.3 ± 29838.7 9534.0 ± 2577.8 0.09 ± 0.04
Tutankham 11.4 291.2 ± 30.4 40.2 ± 14.5 0.10 ± 0.05
Up and Down 533.4 429418.3 ± 7187.4 6072.0 ± 2283.3 0.01 ± 0.01
@@
Video Pinball 0.0 441507.9 ± 283264.6 7943.0 ± 8351.2 0.02 ± 0.02
Wizard of Wor 563.5 49333.3 ± 16157.1 1306.0 ± 1139.8 0.02 ± 0.02
Yars Revenge 3092.9 270262.9 ± 161816.0 8597.4 ± 4291.8 0.02 ± 0.02
Zaxxon 32.5 73097.2 ± 14825.8 896.0 ± 1172.7 0.01 ± 0.02

### v2

@@ Abstract
The search for a general model that can operate seamlessly across multiple domains remains a key goal in machine learning research. The prevailing methodology in Reinforcement Learning (RL) typically limits models to a single task within a unimodal framework, a limitation that contrasts with the broader vision of a versatile, multi-domain model. In this paper, we present Jack of All Trades (JAT), a transformer-based model with a unique design optimized for handling sequential decision-making tasks and multi-modal data types. The JAT model demonstrates its robust capabilities and versatility by achieving strong performance on very different RL benchmarks, along with promising results on Computer Vision (CV) and Natural Language Processing (NLP) tasks, all using a single set of weights. The JAT model marks a significant step towards more general, cross-domain AI model design, and notably, it is the first model of its kind to be fully open-sourced [^1] https://huggingface.co/jat-project/jat, including a pioneering general-purpose dataset.
@@ 1 Introduction
While combining visual and textual tasks has been well-researched, integrating RL tasks remains relatively unexplored and poses distinct challenges. RL tasks are inherently diverse and heterogeneous, making their combination among themselves and with other domains a highly complex exercise. This integration requires dealing with a landscape of different modalities, task complexities, and data volumes across domains and tasks. New questions that arise include: (1) How to design a model and learning method that effectively handles different modalities and data types (sequential decision-making and text-centric)? (2) How to formulate a learning objective that appropriately balances and harmonizes the different modalities, tasks, and domains without bias toward any particular domain or task? (3) How to design a learning strategy that can accommodate the different levels of complexity inherent in different tasks?
These goals are concurrent and, to our knowledge, have only been addressed together by Reed et al. (2022) with the Gato model. Our contributions are characterized by three major advances: (1) Our model features an innovative structure optimized for sequential decision-making tasks. It uniquely assigns each timestep to a corresponding token embedding, resulting in a simpler design. This approach significantly expands the attention window in terms of timesteps compared to Gato (e.g., it is 19 times larger for Atari and more than 25 times larger for Meta-World). (2) In the spirit of open source, we release our code, dataset, and model to the research community. (3) We add observation prediction as an auxiliary task to our model. We demonstrate that this integration significantly contributes to learning a more efficient agent.
@@ 2.1 Transformer for RL
Transformer models (Vaswani et al., 2017) are designed to model sequences and in particular sequences of words in natural language. However, sequence modeling problems span over a much larger set of domains than only NLP. Several efforts have been made to leverage these models for RL (Li et al., 2023). In this paper, we focus on modeling RL trajectories (i.e., sequences of observations, actions and rewards). Modeling such sequences with a Transformer was introduced by Chen et al. (2021) and the Decision Transformer (DT) model. In DT, a Transformer model is trained with offline RL to take sequences of transitions as input and predict the next action. In particular, Chen et al. (2021) proposed to use returns-to-go (i.e. the return from the current state) to condition actions’ generation on both the previous observation and the desired return-to-go. While this has the advantage of explicitly modeling the relations between action selection and return, using the model at inference requires providing at every step a desired return. Liu & Abbeel (2023) proposed to extend this by using hindsight relabelling to better exploit sub-optimal trajectories. Zheng et al. (2022) also extended the DT approach by mixing offine pretraining and online finetuning. Finally, Lee et al. (2022) studied how the DT approach scales to a multi-task RL setup where a single policy is learned for multiple games. In terms of sequence structure, some discretize each dimension of the observation and action spaces separately (Reed et al., 2022; Janner et al., 2021; Chebotar et al., 2023), while others associate an embedding with each element of the sequence (Chen et al., 2021; Zheng et al., 2022).
Our work lies in this line of work as it also leverages Transformers to model trajectories. However, our approach (1) uses standard Behavior Cloning (BC) instead of conditional BC, relaxing the need to condition the agent by the return-to-go and (2) models a multi-task dataset in which sequences come from very different domains (e.g. control, Atari, visual question answering, see Section 3.2).
@@ 3 Methodology
In this section, we introduce the JAT model, detailing our architectural choices that underpin its effectiveness and highlighting its ability to handle different modalities in both sequential decision-making and text-centric tasks. We present the associated dataset, which is notable for its groundbreaking diversity across domains and modalities. Finally, we discuss in depth the learning strategy used.
@@ 3.1.1 Embedding Mechanism
The model is designed to handle two main categories of data: tasks involving sequential decision-making and text-centric tasks. In text-centric tasks, the model currently supports two modalities: text and image. Although the current version of the model supports image generation, we focus only on tasks that involve text generation. To ease reading, we will refer to it as text-centric tasks in the remainder of this paper. Each of these two categories requires a slightly different approach to the embedding process.
@@
For sequential decision-making tasks, the data comprises a sequence of observations, actions, and rewards. At the embedding stage, these sequences are processed to produce an interleaved sequence of observation embeddings (augmented with the corresponding reward) and action embeddings, denoted as $[\phi(s_{0},0.0),\phi(a_{0}),\phi(s_{1},r_{1}),\phi(a_{1}),\ldots]$. Unlike DT (Chen et al., 2021) and Gato (Reed et al., 2022), each timestep is consistently associated with two embeddings: one for the observation and the other for the action, regardless of the modality. This enables JAT to better handle high-dimensional observations, and to provide a much wider, constant attention window in terms of timesteps. As an example, this multiplies the size of the attention window in terms of timesteps by more than 25 for Meta-World. The embedding method employed at a specific timestep is modality-dependent (with $H$ the hidden size of the model):
@@ 3.2.2 Text-Centric Datasets
Refer to caption Refer to caption Refer to caption
@@ 4 Experiments and Results
In this section we discuss the results of our experiments. First, we provide a brief overview of the model’s performance on text-centric tasks. We then present the results of the sequential decision-making tasks, showing the different levels of mastery across the different domains within our study. Finally, we provide a comprehensive analysis highlighting the benefits of incorporating the prediction of the next observation as an auxiliary task during the learning process.
@@ 4.2 Sequential Decision-Making Tasks
The final agent achieves a average expert normalized interquartile mean (IQM) of 65.8%, demonstrating the network’s ability to effectively mimic expert agents across a wide range of tasks. The agent achieves 14.1% of the expert’s score on the Atari 57 benchmark, corresponding to 37.6% of human performance, and exceeding the average human level in 21 games. For the BabyAI benchmark, JAT achieves a normalized score of 99.0%. This score falls below 50% for only one task, namely Move Two Across S8N9. For this benchmark, however, there is no guarantee that the expert score can be achieved, since the bot used for dataset collection has access to the full state of the environment, while the interacting agents only have access to a partial observation. Finally, in the MuJoCo and Meta-World, JAT records scores of 84.8% and 65.5%, respectively. Although JAT reaches expert level for a fair number of Meta-World tasks, we note that some, such as basketball, have not been learned at all. Insofar as the action and observation spaces are identical for all tasks in this benchmark, these failures may be due to task indeterminacy, which we explore in more detail in Appendix D. Future research will have to confirm this hypothesis. We also note that some domains are mastered more quickly than others; in particular, BabyAI achieves a score of 90% after only 30,000 learning steps. We hypothesize that the high semantic similarity of the tasks enables a strong positive transfer, without however providing any proof of this. The Appendix A presents the final results in detail.
@@
Although the results achieved are commendable, for a fair comparison we limit our benchmarking to Gato only, as it is the only truly comparable baseline. Reed et al. (2022) present results only for the 1.18 billion parameter version of Gato, which is 6 times larger than JAT. Its results are normalized to expert performance. Since we don’t have access to the normalization parameters, we estimated scores for the random agents, which may not be exactly the same as those used by Reed et al. (2022), and used our expert scores for normalization, even though they obviously do not match those used by Reed et al. (2022). Therefore, comparisons of these normalized scores should be interpreted with great caution. On the Atari benchmark, JAT achieves an average normalized score of 31.1% outperforming Gato, which reports a score of 30.9%. For BabyAI, JAT achieves an average normalized score of 86.2%, close to the Gato score of 93.2%. Our study, however, is made with 39 tasks versus the 46 used in Gato’s training, with the specific seven additional tasks in their study remaining unidentified. Since our evaluation includes all of the hardest tasks mentioned in their study, the seven missing tasks are likely to be easier, suggesting a harder test scenario in our study. For Meta-World, JAT achieves an average normalized score of 62.8%, which is below the 87.0% reported for Gato. On the MuJoCo benchmark, JAT achieves an average normalized score of 73.6%. While Gato’s training doesn’t use MuJoCo, it’s worth noting that they use the DMC benchmark, which shares some similarities. For reference, on the DMC benchmark (Tassa et al., 2018), Gato achieves an average score of 63.6%.
@@ 5 Conclusion
In this study, we introduce JAT, a novel multi-modal framework for general RL agents. JAT features the ability to handle diverse tasks of varying complexity using a single set of parameters. Its innovations include a new transformer-based structure that efficiently addresses sequential decision-making, CV, and NLP tasks. We also show that joint learning of observation prediction significantly improves performance in sequential decision-making tasks. We’ve open-sourced our training dataset, which includes a wide range of sequential decision-making data as well as extensive language and visual data. We believe that JAT represents an important and valuable step towards general-purpose RL models.
@@
Broader Impact
What sets generalist agents apart is their ability to produce output in a wide range of modalities (textual, visual, virtual or physical control) for multiple applications. This versatility introduces a potential for cross-domain generalisation, although no work to our knowledge actually demonstrates this transfer. The theoretical risk is that such agents transpose behaviours from one domain to another in an inappropriate or undesirable way, and raises important ethical and safety questions. For example, translating aggressive actions in a virtual environment (which may be legitimate in the context of a game or film) into harmful behaviour in the real world. While the alignment of large language models (LLMs) with human values and preferences has been extensively studied (Ouyang et al., 2022; Rafailov et al., 2023; Azar et al., 2023), the application of these alignment strategies to the broader category of generalist models has not been examined in depth. The adaptation of existing alignment methodologies to the requirements of generalist agents is necessary to enable more reliable and secure models.
@@ Acknowledgements
CO2 Emission Related to Experiments
Experiments on the Jean Zay supercomputer (France), with a carbon efficiency of 57.3 gCO2eq/kWh (Statista Search Department, 2020), consumed 2000 GPU hours on Tesla V100-SXM2-32GB hardware (TDP [^5] Estimate for 11/2020: http://www.idris.fr/media/jean-zay/jean-zay-conso-heure-calcul.pdf: 482 W). Total emissions were estimated at 57.2 kgCO2eq, or 29.7 kgCO2eq when considering heat recovery. Estimations used the MachineLearning Impact calculator from (Lacoste et al., 2019).
@@ Appendix A Full results
This appendix contains a detailed view of the results of the trained JAT agent. The score of the random agent for Atari games is sourced from (Mnih et al., 2015). In other domains, this score is approximated by averaging the returns from 1,000 episodes, where the agent selects actions uniformly across its action space. The expert scores represent the average return in the dataset for the task. Meanwhile, the raw score is the average return achieved by the trained agent, based on 100 evaluation episodes. Both these scores, along with the trained agent, are accessible as open-source [^6] https://huggingface.co/jat-project/jat. The normalized score is derived by comparing the agent’s return to the expert’s, calculated using the formula: $\frac{\text{{score}}-\text{{random\_score}}}{\text{{expert\_score}}-\text{random\_score}}$. It’s important to note that in instances where the expert, inaccurately named, does not fully master the task and thus scores similarly or lower than the random agent, the normalized score must be interpreted cautiously. Specifically, if this score falls below that of the random agent, as in the case of Bowling, normalization is not applied. The results for Atari are presented in Table 1 and Figure 7, for BabyAI in Table 2 and Figure 8, for Meta-World in Table 3 and Figure 9, and for MuJoCo in Table 4 and Figure 10.
@@
Alien 227.8 16912.5 ± 7087.4 1474.9 ± 588.7 0.07 ± 0.04
Amidar 5.8 2164.7 ± 1229.5 104.9 ± 103.5 0.05 ± 0.05
Assault 222.4 15699.1 ± 9572.1 1650.1 ± 821.0 0.09 ± 0.05
Asterix 210.0 3699.6 ± 2421.3 800.0 ± 584.9 0.17 ± 0.17
Asteroids 719.0 $1$ 35334.2 1385.3 ± 507.5 0.00 ± 0.00
Atlantis 12850.0 320679.6 ± 418247.4 66980.0 ± 158449.7 0.18 ± 0.51
Bank Heist 14.2 1322.4 ± 60.8 948.3 ± 199.9 0.71 ± 0.15
Battle Zone 236.0 295592.6 ± 161961.0 17420.0 ± 6071.5 0.06 ± 0.02
Beam Rider 363.9 29589.3 ± 16133.0 797.3 ± 328.3 0.01 ± 0.01
Berzerk 123.7 57085.3 ± 13104.5 687.3 ± 331.9 0.01 ± 0.01
Bowling 23.1 20.4 ± 7.3 22.4 ± 5.6 N/A
Boxing 0.1 98.0 ± 3.8 90.1 ± 23.0 0.92 ± 0.24
Breakout 1.7 703.0 ± 203.6 8.8 ± 5.6 0.01 ± 0.01
Centipede 2090.9 11624.3 ± 4918.3 5589.9 ± 2567.3 0.37 ± 0.27
Chopper Command 811.0 90990.6 ± 270876.9 2417.0 ± 1489.9 0.02 ± 0.02
Crazy Climber 10780.5 179296.9 ± 39862.1 97639.0 ± 26184.7 0.52 ± 0.16
Defender 2874.5 351958.3 ± 40466.8 39323.5 ± 15203.0 0.10 ± 0.04
Demon Attack 152.1 92195.2 ± 26174.8 815.3 ± 989.7 0.01 ± 0.01
Double Dunk -18.6 20.9 ± 3.6 14.4 ± 10.0 0.84 ± 0.25
Enduro 0.0 $2$ 147.5 108.5 ± 42.7 0.05 ± 0.02
Fishing Derby -91.7 7.2 ± 25.1 -30.4 ± 24.4 0.62 ± 0.25
Freeway 0.0 33.9 ± 0.3 27.5 ± 1.6 0.81 ± 0.05
Frostbite 65.2 13196.1 ± 4341.0 2769.6 ± 1445.6 0.21 ± 0.11
Gopher 257.6 81676.2 ± 46329.5 5340.6 ± 2547.1 0.06 ± 0.03
Gravitar 173.0 3986.6 ± 1729.0 1269.5 ± 903.0 0.29 ± 0.24
H.E.R.O. 1027.0 44677.4 ± 1754.4 11709.6 ± 3233.5 0.24 ± 0.07
Ice Hockey -11.2 25.2 ± 5.8 7.5 ± 5.6 0.51 ± 0.15
James Bond 29.0 27786.9 ± 33819.2 327.5 ± 123.2 0.01 ± 0.00
Kangaroo 52.0 574.0 ± 636.9 378.0 ± 344.0 0.62 ± 0.66
Krull 1598.0 11439.8 ± 1218.3 10720.5 ± 1284.1 0.93 ± 0.13
Kung-Fu Master 258.5 32392.8 ± 10006.6 288.0 ± 255.1 0.00 ± 0.01
@@
Ms. Pacman 307.3 6896.1 ± 2032.0 1573.1 ± 484.0 0.19 ± 0.07
Name This Game 2292.3 22991.2 ± 2473.1 7523.3 ± 2471.4 0.25 ± 0.12
Phoenix 761.5 424583.2 ± 97649.2 2197.9 ± 1795.4 0.00 ± 0.00
PitFall -229.4 -1.4 ± 4.5 -6.7 ± 19.0 0.98 ± 0.08
Pong -20.7 21.0 ± 0.2 13.7 ± 13.3 0.82 ± 0.32
Private Eye 24.9 100.0 ± 0.0 44.0 ± 49.6 0.25 ± 0.66
Q*Bert 163.9 42971.4 ± 85070.7 1951.5 ± 2577.2 0.04 ± 0.06
River Raid 1338.5 14800.9 ± 7924.6 3758.5 ± 1536.7 0.18 ± 0.11
Road Runner 11.5 77942.8 ± 6088.6 6407.0 ± 4847.4 0.08 ± 0.06
Robotank 2.2 80.5 ± 13.3 11.3 ± 5.5 0.12 ± 0.07
Seaquest 68.4 2597.3 ± 386.1 804.0 ± 403.3 0.29 ± 0.16
Skiing -17098.0 -10738.1 ± $1$16231.5 ± 6060.5 0.14 ± 0.95
Solaris 1236.3 1353.7 ± 517.0 1286.6 ± 446.7 0.43 ± 3.81
Space Invaders 148.0 29425.3 ± 23623.9 325.4 ± 163.4 0.01 ± 0.01
Star Gunner 664.0 360588.6 ± 49207.7 4379.0 ± 3027.2 0.01 ± 0.01
Surround -10.0 9.4 ± 0.8 2.7 ± 4.7 0.65 ± 0.24
Tennis -23.8 11.1 ± 7.6 -13.5 ± 3.8 0.30 ± 0.11
Time Pilot 3568.0 69583.3 ± 29838.7 13028.0 ± 5222.6 0.14 ± 0.08
Tutankham 11.4 291.2 ± 30.4 85.7 ± 61.8 0.27 ± 0.22
Up and Down 533.4 429418.3 ± 7187.4 17768.7 ± 10322.0 0.04 ± 0.02
@@
Video Pinball 0.0 441507.9 ± 283264.6 11917.4 ± 8204.3 0.03 ± 0.02
Wizard of Wor 563.5 49333.3 ± 16157.1 2544.0 ± 2902.4 0.04 ± 0.06
Yars Revenge 3092.9 270262.9 ± 161816.0 12532.7 ± 8062.8 0.04 ± 0.03
Zaxxon 32.5 73097.2 ± 14825.8 6902.0 ± 3206.1 0.09 ± 0.04
Figure 7: Expert normalized episodic return for the JAT agent on the Atari 57 benchmark.